What is Newton second law



1
Expert's answer
2021-07-14T13:47:09-0400

Answer


According to second law of Newton "applied force on a body will be equal to change rate of momentum".

So

"F=\frac{d(P)}{dt}"


Where "F" is force and " P" is momentum.

And

"P=mv"

So

"F=\frac{d(mv)}{dt}"


If mass is constant then

"F=m\frac{d(v)}{dt}"


And "\frac{d(v) }{dt}=a=acceleration"


So

"\colorbox{aqua}{$F=ma$}"